How do you classify a triangle with 8cm 12cm and 17cm side lengths?
Obtuse
Acute
Isosceles
Right

Answer by stanbon(75887)

How do you classify a triangle with 8cm 12cm and 17cm side lengths?
----------
Check c^2 = b^2 + a^2
----------
17^2 = 289
8^2+12^2 = 64+144 = 208
---------------------------
289 > 208
So the angle opposite the 17cm side is more that 90 degrees
Ans: Obtuse
-------------------------
Cheers,
Stan H.
==============
Obtuse
Acute
Isosceles
Right